OpenStudy (anonymous):

the area of a rectangular field is equal to 300 square meters. its perimeter is equal to 70 meters. find the length and width of this rectangle.

OpenStudy (anonymous):

Length of the Rectangular Field is 20 m and Width is 15 m: Let L: Length of the Rectangle W: Width of Rectangle Area of Rectangular Field : Length x Width = LxW =300---(i) Perimeter = 2(L+W)=70 (L+W)=35 L=35-W Put this value in Eq (i), you will get following equation : x2-35x+300 =0 After solving this equation, L=20 and W=15

OpenStudy (anonymous):